OpenStudy (anonymous):

What are the solutions of 6x2 = 12x? Please explain to me how to do this, and not just the answer thx :)

OpenStudy (aaronq):

you can make it equal to zero, 6x^2-12x=0 then factor

OpenStudy (anonymous):

6x^2-12x=0 +12x 12x 6x^2=12x now what

OpenStudy (aaronq):

you reverted back to the original format given. you need to factor out the x 6x^2-12x=0 x(6x-12)=0 then x=0, and 6x-12=0 -> x=12/6=2

OpenStudy (anonymous):

do I plug x back in?

OpenStudy (aaronq):

nope, you're trying to find what the values (roots) of x are. The roots are where the graph crosses the x axis.

OpenStudy (anonymous):

x = 0, x = −2 x = 0, x = 2 x = 6, x = 2 x = 6, x = −2

OpenStudy (anonymous):

gosh im not getting this at all

OpenStudy (aaronq):

dont stress and think about it. you have: 6x^2=12x you need to make it equal to zero to find the values of x. 6x^2-12x=0 then you factor the x out, x(6x-12)=0 so x=0 and 6x-12=0 just solve second part from the bracket and thats your second root.

OpenStudy (anonymous):

6x^3-12x=0

OpenStudy (anonymous):

becuase x*6x=6x^2 x*-12=-12x

OpenStudy (aaronq):

6x^3=12x is the actual question?

OpenStudy (anonymous):

the question is [9.02] What are the solutions of 6x2 = 12x? x = 0, x = −2 x = 0, x = 2 x = 6, x = 2 x = 6, x = −2

OpenStudy (aaronq):

okay, so as i said, make it equal to zero: 6x^2=12x subtract 12x from both sides 6x^2-12x=0 factor out 1 of the x's (because you can't factor both) you can also factor out 6, 6x(x-2)=0 so you have: 6x=0 and x-2=0

OpenStudy (aaronq):

now you need to solve for x (by isolating it)

OpenStudy (anonymous):

I thought you have to do the opposite of what the sign tells you, the -12 so I would have to add from both sides to cancel out

OpenStudy (aaronq):

well yes, if you have x-12=0 to isolate x, you add 12 to both sides x-12+12=12 x=12 is that what you're getting at?

OpenStudy (anonymous):

how to I give you a medal or something on here for being so patient (new to this site) sorry but could you please sjow me how to solve it all together, with step 1 , 2 ect

OpenStudy (aaronq):

no problem, if i wasn't patient i wouldn't be helping ya lol you have: 6x^2=12x you wanna make it equal to zero to solve for x. bring the 12x to the other side, since it's positive it's adding notice that if you wanted to isolate the x, like this 6x^2/12=x then yes you would divide. but thats not what were doing. 6x^2-12x=12x-12x -> 6x^2-12x=0 makes sense?

OpenStudy (anonymous):

yes, ok so you made it so it would equal 0

OpenStudy (aaronq):

now we factor what we can 6x^2-12x=0 factor the x from both 6x^2 and -12x x(6x-12)=0 we can also factor the 6 6x(x-2)=0 now we can separate them since both parts when multiplied equal zero 6x=0 and x-2=0 now it's just a matter of isolating x

OpenStudy (anonymous):

I isolate what? :O

OpenStudy (aaronq):

isolate x from "6x=0" and "x-2=0"

OpenStudy (anonymous):

so x=6 and x=2

OpenStudy (aaronq):

6x=0 divide both sides by 6 6x/6=0 -> x=0

OpenStudy (aaronq):

no, here 6x is multiplied, you need to divide both sides by 6

OpenStudy (anonymous):

6/x

OpenStudy (aaronq):

nononono 6x=0 -> 6x/6=0/6 right?

OpenStudy (aaronq):

you're doing it as if it was x-6=0, which is not the case

OpenStudy (anonymous):

I'm about to jump out a window ... jk but for real I am sooo lost it seems so simple but math is so confusing D;

OpenStudy (aaronq):

you're getting confused with the algebraic operations of when to divide/multiply and when to subtract/add. haha id suggest you review those concepts

OpenStudy (anonymous):

ok so , the answer is 6x=0 -> 6x/6=0/6

OpenStudy (anonymous):

Nvm I feel bad, haha I dont want to waste youre time any more

OpenStudy (aaronq):

lol don't worry about it. yes, the roots are 6x/6=0/6 -> x=0 and x=2

OpenStudy (aaronq):

do you understand it? or what do you not understand?

OpenStudy (anonymous):

after that what do I need to do?

OpenStudy (aaronq):

thats it x=0 and x=2 are the roots, which is what you were looking for

OpenStudy (anonymous):

I feel so dumb, ok thank you for your help!!! you rox :)

OpenStudy (aaronq):

haha you're not dumb. all you need is practice, so practice some more! :)